OpenStudy (anonymous):

Using complete sentences, explain how to find the equation of the line, in standard form and slope–intercept form, passing through (–1, 5) and (1, 9). Also compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form. QUICK! and Simple

OpenStudy (anonymous):

X-X1/X1-X2=Y-Y1/Y1-Y2 Y=2X+7

OpenStudy (anonymous):

Complete sentences please:(